Circular No. 6791

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 971214 A. Diercks and E. W. Deutsch, University of Washington; R. Wyse,

The Johns Hopkins University; G. Gilmore, Cambridge University; C. Corson, Apache Point Observatory; F. J. Castander, University of Chicago; and E. Turner, Princeton University, write: “We confirm the fading optical transient reported by Halpern et al. on IAUC 6788, which we measure to be at α = 11^{h}56^{m}26^{s}.35, δ = +65^{o}12′00′′.7 (equinox 2000.0). The object was observed to fade by ∼ 1.6 mag in R over 24 hr in images taken by Corson, Gilmore, and Wyse with the ARC 3.5-m telescope at Apache Point: Dec. 15.51 UT, R = 22.1 ± 0.1; 16.52, 23.7 ± 0.3. No other object brighter than R ∼ 21.5 was observed to vary. Magnitudes and positions were calculated based on photometry and astrometry of two stars measured in the ‘Automatic Plate Measuring’ (APM) survey (Maddox et al. 1990, MNRAS 243, 692); neither star varied (within uncertainties) over the 24-hr period, their position end figures and magnitudes being 25^{s}.75, 11′35′′.7, R = 20.10 ± 0.01; 34^{s}.22, 11′44′′.2, 20.78 ± 0.03. Estimated zeropoint errors in converting from the APM R_{F} band to the R band are ∼ 0.3 mag. Our R-band observations on Dec. 15 and 16 were approximately coincident with the reported MDM I-band detections by Halpern et al., giving a color for this object of R − I = +1.0 ± 0.4. Within the quoted errors, the color is constant during the fading. Images are available at

http://www.astro.washington.edu/deutsch/misc/grb971214/.”

Castander, Wyse, Gilmore, Corson, Diercks, and Deutsch; D. Q. Lamb, University of Chicago; and Turner report an additional magnitude from six 10-min exposures (using Landolt standard star 104239), indicating that the observed R-band fluxes are consistent with a power-law decay in time: Dec. 17.51 UT, R = 24.4 ± 0.5. The first of the unvarying stars noted above was measured at R = 20.14 ± 0.02 at this time.

Regarding their item on IAUC 6788, N. Itoh et al. note that their faint object is near HD 103690; though the position accuracy is limited by poor seeing, they provide the following improved position end figures: 25^{s}.6 ± 0^{s}.2, 55′′.7 ± 0′′.1. Also, on line 1, for done read down

COMET 103P/HARTLEY 2 Total visual magnitude estimates by M. Lehký, Hradec Králové, Czech

Republic (L = 0.42-m reflector; B = 25×100 binoculars): Oct. 19.76 UT, 11.1 (L); 21.75, 11.6 (L); 31.84, 10.9 (L); Nov. 4.74, 10.7 (L); 10.73, 10.7 (L); 20.71, 9.7 (B); Dec. 14.74, 9.1 (B); 16.72, 8.4 (B).

1997 December 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green